# Polarity transitions induced by symmetry-breaking outer boundary heat flux in rapidly rotating dynamos

### v1

@@ 3.3 The role of slow MAC waves in the dipole–multipole transition
Figure 13: (a-b) Contour plots of $\partial{u}_{z}/\partial t$ at a cylindrical radius of $s=1$ are shown for the scales $l\leq l_{E}$ over short time intervals in the saturated state of two dynamo simulations with $q^{\star}=17$ and $18$. The parallel black lines represent the primary wave travel direction, with their slope giving the group velocity $U_{g,z}$. Table 3 lists the estimated group velocities for the fast and slow MAC waves ($U_{f}$ and $U_{s}$, respectively) and $U_{g,z}$. The other dynamo parameters are $Ra_{V}=2500,E=1.2\times 10^{-5}$, and $Pm=Pr=1$.
SL $E$ $Ra_{V}$ $q^{\star}$ Fig. $\omega_{n}^{2}$ $\omega_{C}^{2}$ $\omega_{M}^{2}$ $-{\omega_{A}}^{2}$ $\omega_{f}$ $\omega_{s}$ $U_{f}$ $U_{s}$ $U_{g,z}$
No. (×$(\times 10^{10})$ (×$(\times 10^{8})$ (×$(\times 10^{8})$ (×$(\times 10^{8})$ (×$(\times 10^{4})$ (×$(\times 10^{4})$
1 $1.2\times 10^{-5}$ 2500 17 13 (a) 2.5 13.02 6.59 3.94 4.63 0.90 8989 525 456
2 $1.2\times 10^{-5}$ 2500 18 13 (b) 2.7 11.51 2.22 4.32 3.46 0.0 9651 0 9274
Table 3: Summary of the data for MAC wave measurement in the dynamo models. The sampling frequency $\omega_{n}$ is selected to ensure that the fast MAC waves are captured when measuring group velocity. The values of $\omega_{M}^{2}$, $-\omega_{A}^{2}$, and $\omega_{C}^{2}$ are computed using (49), based on the averaged wavenumbers $m$, $k_{s}$, and $k_{z}$ in the energy-containing scales where $l\leq l_{E}$. The group velocity in the $z$ direction ($U_{g,z}$) is then compared with the estimated velocities of the fast ($U_{f}$) and slow ($U_{s}$) MAC waves.
Figure 13 shows the measurement of wave velocities from their propagation paths in the saturated state of dynamos at $q^{\star}=17$ and $q^{\star}=18$. Contours of the fluctuating $z$-velocity, represented by $\partial u_{z}/\partial t$ at cylindrical radius $s=1$, are plotted over short time intervals where the ambient magnetic field and wavenumbers in the energy-containing scales $l\leq l_{E}$ remain approximately constant. By examining the slope of the black lines, we determine the axial group velocity $U_{g,z}$ and compare it with the estimated axial group velocities of the fast ($U_{f}$) and slow ($U_{s}$) waves. These velocities are derived from the respective frequencies in the diffusionless limit by taking their derivative with respect to $k_{z}$ (Varma & Sreenivasan, 2022). In the dipole-dominated run at $q^{\star}=17$, the the slow MAC waves are predominant although the fast MAC waves also exist (figure 13(a)). In contrast, at $q^{\star}=18$, the slow MAC waves are absent while the fast waves are abundant (figure 13(b)).
@@ 3.4 Complementarity of vertical and horizontal buoyancies in the dipole–multipole transition
Figure 14: (a) Variation of the modified vertical Rayleigh number $Ra_{V}$ with the peak Elsasser number $\varLambda$ (square of the peak magnetic field) at the suppression of slow MAC waves. (b) Variation of the local Rayleigh number $Ra_{\ell}$, defined in (51), with $\varLambda$. The values of $Ra_{V}$, $Ra_{\ell}$ and $\varLambda$ in the plots are given in table 2. The parameters of the two dynamo series and their symbolic representations are as follows: $E=6\times 10^{-5},Pm=Pr=5$ (circles), $E=1.2\times 10^{-5},Pm=Pr=1$ (diamonds). The filled symbols show the evolution path of the dynamo with increasing heat flux heterogeneity. Black symbols represent the dipolar state while red symbols represent the reversing or multipolar state. (c) The states of polarity reversals are shown in a plot of local relative vertical versus relative horizontal Rayleigh numbers, normalized by the resultant local Rayleigh number $Ra_{\ell}$. For $E=1.2\times 10^{-5}$, reversals occur at $Ra_{V}=27000$ and $Ra_{\ell,H}/Ra_{\ell}=0.2$ (blue diamond) as well as at $Ra_{V}=1500$ and $Ra_{\ell,H}/Ra_{\ell}=0.64$ (red diamond). (d) Variation of $q^{\star}$ with respect to $Ra_{\ell,H}/Ra_{\ell}$. The horizontal line at $q^{\star}=10$ and the vertical line at $Ra_{\ell,H}/Ra_{\ell}=0.5$ indicate that $Ra_{\ell,H}/Ra_{\ell}>0.5$ corresponds to $q^{\star}=O(10)$.
@@
In the presence of rapid rotation and a magnetic field, the buoyant forcing generates fast and slow MAC waves when the inequality $|\omega_{C}|>|\omega_{M}|>|\omega_{A}|>|\omega_{\eta}|$ is satisfied. The magnetic field intensity increases with the vigour of convection in this dipole-dominated regime; however, a sufficiently strong forcing suppresses the slow MAC waves when $|\omega_{A}|\approx|\omega_{M}|$, leading to dipole collapse. The nature of buoyancy could be either vertical, measured by $Ra_{V}$, or horizontal, measured by $Ra_{H}$. In a dynamo with homogeneous boundary heat flux, the progressive increase of $Ra_{V}$ produces reversals in a narrow range of forcing between the dipolar and multipolar regimes (Majumder et al., 2024). In this study, $Ra_{H}$ is progressively increased for a fixed $Ra_{V}$ until the dynamo undergoes the polarity transition (table 1). The Elsasser number at the polarity transition, $\varLambda=B^{2}_{\text{peak}}$ is shown against $Ra_{V}$ for two different Ekman numbers in figure 14(a). The resultant local Rayleigh number $Ra_{\ell}$, defined by (52)), on the other hand, varies in self-similar fashion with $\varLambda$ (figure 14(b)). This self-similar line separates the dipolar and multipolar regimes. The filled symbols represent the evolution path of the dynamo with increasing order of heat flux heterogeneity for two vertical Rayleigh numbers ($Ra_{V}=2500$ and $Ra_{V}=20000$). The filled black symbols here represent the states where $|\omega_{M}|>|\omega_{A}|$ while the filled red symbols represent $|\omega_{M}|\lesssim|\omega_{A}|$. In figure 14(c), all the polarity-reversing states are plotted in the $Ra_{\ell,H}$ versus $Ra_{\ell,V}$ space, normalized by the value of $Ra_{\ell}$. This plot also demarcates the dipolar and multipolar regimes and furthermore demonstrates the complementarity between the horizontal and vertical buoyancies in producing the polarity transition through the condition of vanishing slow MAC waves, as suggested by figure 3(e). For $E=1.2\times 10^{-5}$, reversals occur in a strongly driven dynamo at $Ra_{V}=27000$ and $Ra_{\ell,H}/Ra_{\ell}=0.2$ (blue diamond) as well as a relatively weakly driven dynamo at $Ra_{V}=1500$ and $Ra_{\ell,H}/Ra_{\ell}=0.64$ (red diamond); see also table 2. Figure 14(d) relates the two measures of heterogeneity, $q^{*}$ and $Ra_{\ell,H}/Ra_{\ell}$ and indicates that values of $Ra_{\ell,H}/Ra_{\ell}>0.5$ correspond to values of $q^{*}$ of $O(10)$.
@@ 3.5 Combination of symmetric and anti-symmetric boundary heat flux
Since the heat flux heterogeneity in the lower mantle is a combination of both equatorially symmetric and anti-symmetric variations, it is instructive to consider the effect of composite heat flux boundary conditions on polarity transitions. The analysis of the heat flux distribution derived from the seismic shear wave velocity in the Earth’s lower mantle (Masters et al., 1996) suggests a complex heat flux pattern at the CMB which features a well-defined $Y_{2}^{2}$ component. That said, the analysis of peak-to-peak heat flux variations reveals that the ratio of the symmetric to anti-symmetric heat flux variation is approximately 1.65.
With a composite heat flux heterogeneity at the outer boundary, a dipole-dominated solution is observed when the inequality $|\omega_{C}|>|\omega_{M}|>|\omega_{A}|$ is satisfied in the presence of the slow MAC waves. For a given vertical buoyant forcing ($Ra_{V}\sim 10^{2}$ times its value for nonmagnetic onset of convection), the progressive increase of horizontal buoyancy produces polarity transitions for composite patterns where the equatorially symmetric and anti-symmetric variations are comparable. For example, a heat flux pattern with $Y_{2}^{2}:Y_{2}^{1}=2:1$ and the Earth-like heat flux pattern derived from seismic tomography both produce reversals for $q^{*}$ of $O(10)$ (table 1). For the same $Ra_{V}$, the values of $Ra_{\ell,H}/Ra_{\ell}$ at which the transition occurs in the two cases are nearly equal, which points to the comparable ratios of the symmetric to anti-symmetric heat flux variation. Notably, a heterogeneity consisting of comparable magnitudes of symmetric and anti-symmetric variations induces the polarity transition at a value of $Ra_{\ell,H}/Ra_{\ell}$ (and $q^{\star}$) which is of the same order as that for the transition induced by a purely anti-symmetric variation.
The polarity transition with the boundary heat flux derived from seismic tomography is analysed further in figure 17 of Appendix A. In the dipole-dominated state at $q^{*}=10$, $|\omega_{M}|^{2}>|\omega_{A}|^{2}$ in several regions, where $\omega_{A}^{2}$ is based on the resultant basic state temperature gradient $\beta$, defined in (46). In these regions, the Alfvén wave velocity is greater than the velocity of the buoyancy perturbations and the slow MAC waves are generated. On the other hand, the polarity-reversing state at $q^{*}=13$ shows approximate parity between $|\omega_{M}|^{2}$ and $|\omega_{A}|^{2}$ so that the slow waves are suppressed. This analysis also explains why a composite boundary heterogeneity made up of a dominant equatorially symmetric variation does not induce polarity transitions. For example, the heterogeneity with the ratio $Y_{2}^{2}:Y_{2}^{1}=5:1$ does not admit polarity transitions even at high $q^{\star}$ (table 1) since the condition $|\omega_{A}|^{2}\gtrsim|\omega_{M}|^{2}$ is not met.
@@ 4 Two-component magnetoconvection
4.1 Evolution of a density disturbance in two-component convection
@@
$\frac{\partial\bm{b}}{\partial t}+u_{c}\,z\frac{\partial\bm{b}}{\partial x}=u_{c}b_{z}\hat{\bm{e}}_{x}+(\bm{B}_{0}\cdot\nabla)\bm{u}+\nabla^{2}\bm{b},$ (56)
@@
The dimensionless parameters in (55), based on the length scale of the density disturbance, are the Ekman number $E=\nu/2\varOmega\delta^{2}$, magnetic Ekman number $E_{\eta}=\eta/2\varOmega\delta^{2}$, and local vertical thermal and compositional Rayleigh numbers, $Ra_{\ell,V}^{T}=g\alpha^{T}|\beta_{y}^{T}|\delta^{2}/2\varOmega\eta$ and $Ra_{\ell,V}^{C}=g\alpha^{C}|\beta_{y}^{C}|\delta^{2}/2\varOmega\eta$, respectively. Here, $\alpha^{T}$ and $\alpha^{C}$ are the thermal and compositional expansion coefficients, respectively. Additionally, $q^{T}=\kappa^{T}/\eta$, $q^{C}=\kappa^{C}/\eta$ and $\beta^{\star}=\beta_{z}/\beta_{y}^{C}$, where $\kappa_{T}$ and $\kappa_{C}$ are the diffusivities of temperature and composition and $\beta_{z}$ is the horizontal temperature gradient.
@@
[…]
@@
$\lambda_{1,2}$$\approx$ ±$\pm\bigg{(}\omega_{C}+\frac{\omega_{M}^{2}}{\omega_{C}}\bigg{)}+\mathrm{i}\,\frac{\omega_{M}^{2}\omega_{\eta}}{\omega_{C}^{2}},$ (62)
$\lambda_{3,4}$$\approx$ ±$\pm\bigg{(}\frac{\omega_{M}^{2}}{\omega_{C}}+\frac{{\omega_{A,V}^{C}}^{2}+{\omega_{A,V}^{T}}^{2}-2\omega_{A,H}^{2}}{2\omega_{C}}\bigg{)}+\mathrm{i}\,\omega_{\eta}\,\bigg{(}1-\frac{{\omega_{A,V}^{C}}^{2}+{\omega_{A,V}^{T}}^{2}-\omega_{A,H}^{2}}{2\omega_{M}^{2}}\bigg{)},$ (63)
@@
$Ra_{\ell}=\frac{g\big{(}\alpha^{C}\beta_{y}^{C}+\alpha^{T}\beta_{y}^{T}-2\alpha^{T}\beta_{z}k_{y}/k_{z}\big{)}\delta^{2}}{2\varOmega\eta},$ (66)
@@
The coefficients $D_{m}$, $G_{m}$ and $Q_{m}$ are evaluated from the initial conditions for $\hat{u}_{z}$, $\hat{\Theta}$, $\hat{\gamma}$ and their time derivatives. For instance, the initial conditions for $\hat{u}_{z}$ and its time derivatives are given by
@@
Further,
$\hat{u}_{y}=-\frac{k_{z}}{k_{y}}\hat{u}_{z}.$ (72)
In a similar way, the spectral coefficients of $\hat{\Theta}$ and $\hat{\gamma}$ are obtained (Appendix B).
For the dipole-dominated regime given by $|\omega_{C}|>|\omega_{M}|>|\omega_{A}|$, the Elsasser number $\varLambda$ based on the peak magnetic field should take values of $O(10^{2})$ (Sreenivasan & Maurya, 2021). Furthermore, if strong compositional buoyancy in the planetary core by itself maintains the dynamo in a dipolar state below the threshold for polarity transitions, a value of $Ra^{C}_{\ell,V}$ of $O(10^{3})$ is reasonable (Majumder et al., 2024). For Earth, this value of the local Rayleigh number corresponds to a vertical Rayleigh number $O(10^{3})$ times the critical Rayleigh number for onset of nonmagnetic convection, essential for reproducing the observed polar circulation of 0.6–0.9 yr-1 (Hulot et al., 2002) in the low-inertia geodynamo (see Majumder & Sreenivasan, 2023). Now, using the complementarity of the buoyancy frequencies (65), the value of the relative horizontal buoyancy $Ra_{\ell,H}/Ra^{T}_{\ell}$ for polarity transitions based on the condition of vanishing slow MAC waves ($|\omega_{A}|\approx|\omega_{M}|$) may be obtained for a range of thermal power fractions $f^{T}$, given by

### v2

@@ 3.3 The role of slow MAC waves in the dipole–multipole transition
Figure 14: (a-b) Contour plots of $\partial{u}_{z}/\partial t$ at a cylindrical radius of $s=1$ are shown for the scales $l\leq l_{E}$ over short time intervals in the saturated state of two dynamo simulations with (a) $q^{*}=17$ and (b) $q^{*}=18$ for the equatorially anti-symmetric $Y_{2}^{1}$ heat flux boundary condition. The parallel black lines represent the primary wave travel direction, with their slope giving the group velocity $U_{g,z}$. Table 3 lists the estimated group velocities for the fast and slow MAC waves ($U_{f}$ and $U_{s}$, respectively) and $U_{g,z}$. The other dynamo parameters are $Ra_{V}=2500,E=1.2\times 10^{-5}$, and $Pm=Pr=1$.
SL $E$ $Ra_{V}$ $q^{*}$ Fig. $\omega_{n}^{2}$ $\omega_{C}^{2}$ $\omega_{M}^{2}$ ${\omega_{A}}^{2}$ $\omega_{f}$ $\omega_{s}$ $U_{f}$ $U_{s}$ $U_{g,z}$
No. $(\times 10^{10})$ $(\times 10^{8})$ $(\times 10^{8})$ $(\times 10^{8})$ $(\times 10^{4})$ $(\times 10^{4})$
1 $1.2\times 10^{-5}$ 2500 17 14 (a) 2.5 13.02 6.59 3.94 4.63 0.90 $89$ 525 456
2 $1.2\times 10^{-5}$ 2500 18 14 (b) 2.7 11.51 2.22 4.32 3.46 0.0 9651 0 9274
Table 3: Summary of the data for MAC wave measurement in the dynamo models. The sampling frequency $\omega_{n}$ is selected to ensure that the fast MAC waves are captured when measuring group velocity. The values of $\omega_{M}^{2}$, $\omega_{A}^{2}$, and $\omega_{C}^{2}$ are computed using (52), based on the averaged wavenumbers $m$, $k_{s}$, and $k_{z}$ in the energy-containing scales where $l\leq l_{E}$. The group velocity in the $z$ direction ($U_{g,z}$) is then compared with the estimated velocities of the fast ($U_{f}$) and slow ($U_{s}$) MAC waves.
Figure 14 shows the measurement of wave velocities from their propagation paths in the saturated state of dynamos at $q^{*}=17$ and $q^{*}=18$ for the equatorially anti-symmetric $Y_{2}^{1}$ heat flux pattern at the outer boundary. Contours of the fluctuating $z$-velocity, represented by $\partial u_{z}/\partial t$ at cylindrical radius $s=1$, are plotted over short time intervals where the ambient magnetic field and wavenumbers in the energy-containing scales $l\leq l_{E}$ remain approximately constant. By examining the slope of the black lines, we determine the axial group velocity $U_{g,z}$ and compare it with the estimated axial group velocities of the fast ($U_{f}$) and slow ($U_{s}$) waves. These velocities are derived from the respective frequencies in the diffusionless limit by taking their derivative with respect to $k_{z}$ (Varma and Sreenivasan, 2022). In the dipole-dominated run at $q^{*}=17$, the slow MAC waves are predominant although the fast MAC waves also exist (figure 14(a)). In contrast, at $q^{*}=18$, the slow MAC waves are nearly absent while the fast waves are abundant (figure 14(b)).
@@ 3.4 Complementarity of vertical and horizontal buoyancies in the dipole–multipole transition
Figure 15: (a) Variation of the modified vertical Rayleigh number $Ra_{V}$ with the square of the peak magnetic field, $B^{2}_{peak}$ at the suppression of slow MAC waves. (b) Variation of the local Rayleigh number $Ra_{\ell}$, defined in (55), with $B^{2}_{peak}$. The values of $_{V}, RaRa_{\ell}$ and $B^{2}_{peak}$ in the plots are given in table 2. The parameters of the two dynamo series and their symbolic representations are as follows: $E=6\times 10^{-5},Pm=Pr=5$ (circles), $E=1.2\times 10^{-5},Pm=Pr=1$ (diamonds). The filled symbols show the evolution path of the dynamo with increasing heat flux heterogeneity. Black symbols represent the dipolar state and red symbols represent the reversing or multipolar state. (c) The states of polarity reversals are shown in a plot of local relative vertical versus relative horizontal Rayleigh numbers, normalized by the resultant local Rayleigh number $Ra_{\ell}$. For $E=1.2\times 10^{-5}$, reversals occur at $Ra_{V}=27000$ and $Ra_{\ell,H}/Ra_{\ell}=0.2$ (blue diamond) as well as at $Ra_{V}=1500$ and $Ra_{\ell,H}/Ra_{\ell}=0.64$ (red diamond). (d) Variation of $q^{*}$ with respect to $Ra_{\ell,H}/Ra_{\ell}$. The horizontal line at $q^{*}=10$ and the vertical line at $Ra_{\ell,H}/Ra_{\ell}=0.5$ indicate that $Ra_{\ell,H}/Ra_{\ell}>0.5$ corresponds to $q^{*}=O(10)$.
@@
In the presence of rapid rotation and a magnetic field, the buoyant forcing generates fast and slow MAC waves when the inequality $|\omega_{C}|>|\omega_{M}|>|\omega_{A}|>|\omega_{\eta}|$ is satisfied. The magnetic field intensity increases with the vigour of convection in this dipole-dominated regime; however, a sufficiently strong forcing suppresses the slow MAC waves when $|\omega_{A}|\approx|\omega_{M}|$, leading to dipole collapse. The nature of buoyancy could be either vertical, measured by $Ra_{V}$, or horizontal, measured by $Ra_{H}$. In a dynamo with homogeneous boundary heat flux, the dipolar regime transitions in succession to polarity-reversing and multipolar states with increasing $Ra_{V}$ (Majumder et al., 2024). In this study, $Ra_{H}$ is progressively increased for a fixed $Ra_{V}$ until the dynamo undergoes the polarity transition (table 1). The square of the peak magnetic field measured at the polarity transition, $B^{2}_{peak}$, is plotted against $Ra_{V}$ for two different Ekman numbers in figure 15(a). The approximately linear variation of $Ra_{V}$ with $B^{2}_{peak}$ is different in the two cases. On the other hand, the resultant local Rayleigh number $Ra_{\ell}$, defined by equation (56), exhibits a self-similar variation with $B^{2}_{peak}$ (figure 15(b)). This self-similar line separates the dipolar and multipolar regimes. The filled symbols represent the evolution path of the dynamo with increasing order of heat flux heterogeneity for two vertical Rayleigh numbers ($Ra_{V}=2500$ and $Ra_{V}=20000$). The filled black symbols here represent the states where $|\omega_{M}|>|\omega_{A}|$ while the filled red symbols represent $|\omega_{M}|\lesssim|\omega_{A}|$. In figure 15(c), all the polarity-reversing states are plotted in the $Ra_{\ell,H}$ versus $Ra_{\ell,V}$ space, normalized by the value of $Ra_{\ell}$. This plot also demarcates the dipolar and multipolar regimes and furthermore demonstrates the complementarity between the horizontal and vertical buoyancies in producing the polarity transition through the condition of vanishing slow MAC waves, as suggested by figure 3(e). For $E=1.2\times 10^{-5}$, reversals occur in a strongly driven dynamo at $Ra_{V}=27000$ and $Ra_{\ell,H}/Ra_{\ell}=0.2$ (blue diamond) as well as a relatively weakly driven dynamo at $Ra_{V}=1500$ and $Ra_{\ell,H}/Ra_{\ell}=0.64$ (red diamond); see also table 2. Since the self-generated magnetic field depends on the vertical Rayleigh number $Ra_{\ell,V}$, only weak-field dynamos of $B^{2}_{rms}<O(1)$ exist for $Ra_{\ell,V}/Ra_{\ell}<0.25$. Therefore, solutions for $Ra_{\ell,V}/Ra_{\ell}<0.25$ (i.e. $Ra_{\ell,H}/Ra_{\ell}>0.75$) are not considered. Figure 15(d) relates the two measures of heterogeneity, $q^{*}$ and $Ra_{\ell,H}/Ra_{\ell}$ and indicates that values of $Ra_{\ell,H}/Ra_{\ell}>0.5$ correspond to values of $q^{*}$ of $O(10)$.
@@ 3.5 Combination of symmetric and anti-symmetric boundary heat flux
Since the heat flux heterogeneity in the lower mantle is a combination of both equatorially symmetric and anti-symmetric variations, it is instructive to consider the effect of composite heat flux boundary conditions on polarity transitions. The analysis of the heat flux distribution based on the seismic shear wave velocity in the Earth’s lower mantle (Masters et al., 1996) suggests a complex heat flux pattern at the CMB which features a well-defined $Y_{2}^{2}$ component. That said, the analysis of peak-to-peak heat flux variations reveals that the ratio of the symmetric to anti-symmetric heat flux variation is approximately 1.65. This ratio is obtained by calculating the peak-to-peak variation in heat flux for the symmetric and anti-symmetric contributions separately, within the seismic tomography pattern. With a composite heat flux heterogeneity at the outer boundary, a dipole-dominated solution is observed when the inequality $|\omega_{C}|>|\omega_{M}|>|\omega_{A}|$ is satisfied in the presence of the slow MAC waves. For a given vertical buoyant forcing ($Ra_{V}\sim 10^{2}$ times its value for nonmagnetic onset of convection), the progressive increase of horizontal buoyancy produces polarity transitions for composite patterns where the equatorially symmetric and anti-symmetric variations are comparable. For example, a heat flux pattern with $Y_{2}^{2}:Y_{2}^{1}=2:1$ and the Earth-like heat flux pattern derived from seismic tomography both produce reversals for $q^{*}$ of $O(10)$ (table 1). For the same $Ra_{V}$, the values of $Ra_{\ell,H}/Ra_{\ell}$ at which the transition occurs in the two cases are nearly equal, which points to the comparable ratios of the symmetric to anti-symmetric heat flux variation. Notably, a heterogeneity consisting of comparable magnitudes of symmetric and anti-symmetric variations induces the polarity transition at a value of $Ra_{\ell,H}/Ra_{\ell}$ (and $q^{*}$) which is of the same order as that for the transition induced by a purely anti-symmetric variation.
The polarity transition with the boundary heat flux derived from seismic tomography is analysed further in figure 18 of Appendix A. In the dipole-dominated state at $q^{*}=10$, $|\omega_{M}|^{2}>|\omega_{A}|^{2}$ in several regions, where $\omega_{A}^{2}$ is based on the resultant basic state temperature gradient $\beta$, defined in (49). In these regions, the Alfvén wave velocity is greater than the velocity of the buoyancy perturbations and the slow MAC waves are generated. On the other hand, the polarity-reversing state at $q^{*}=13$ shows approximate parity between $|\omega_{M}|^{2}$ and $|\omega_{A}|^{2}$ so that the slow waves are suppressed. This analysis also explains why a composite boundary heterogeneity made up of a dominant equatorially symmetric variation does not induce polarity transitions. For example, the heterogeneity with the ratio $Y_{2}^{2}:Y_{2}^{1}=5:1$ does not admit polarity transitions even at high $q^{*}$ (table 1) since the condition $|\omega_{A}|^{2}\gtrsim|\omega_{M}|^{2}$ is not met.
@@ 4 Two-component magnetoconvection
4.1 Evolution of a density disturbance in two-component magnetoconvection
@@
The dimensionless parameters in (58), based on the length scale of the density disturbance, are the Ekman number $E=\nu/2\varOmega\delta^{2}$, magnetic Ekman number $E_{\eta}=\eta/2\varOmega\delta^{2}$, and local vertical compositional and thermal Rayleigh numbers, $Ra_{\ell,V}^{C}=g\alpha^{C}|\beta_{y}^{C}|\delta^{2}/2\varOmega\eta$ and $Ra_{\ell,V}^{T}=g\alpha^{T}|\beta_{y}^{T}|\delta^{2}/2\varOmega\eta$, respectively. Here, $\alpha^{T}$ and $\alpha^{C}$ are the thermal and compositional expansion coefficients, respectively. Additionally, $q^{T}=\kappa^{T}/\eta$, $q^{C}=\kappa^{C}/\eta$ and $\beta^{\star}=\beta_{z}/\beta_{y}^{C}$, where $\kappa_{T}$ and $\kappa_{C}$ are the diffusivities of temperature and composition and $\beta_{z}$ is the horizontal temperature gradient.
@@
$\begin{aligned} &\omega_{C}^{2}=\frac{1}{E_{\eta}^{2}}\frac{k_{z}^{2}}{k^{2}},\quad\omega_{M}^{2}=\frac{(\bm{B}\cdot\bm{k})^{2}}{E_{\eta}},\quad{\omega_{A,V}^{T}}^{2}=\frac{Ra_{\ell,V}^{T}}{E_{\eta}}\frac{k_{z}^{2}}{k^{2}},\quad{\omega_{A,V}^{C}}^{2}=\frac{Ra_{\ell,V}^{C}}{E_{\eta}}\frac{k_{z}^{2}}{k^{2}},\quad\\ &\omega_{A,H}^{2}=\frac{Ra_{\ell,H}}{2E_{\eta}}\frac{k_{z}^{2}}{k^{2}},\quad\omega_{\eta}^{2}=k^{4}.\end{aligned}$ (64a–f)
@@
$\lambda_{1,2}$ $\approx$ $\pm\bigg(\omega_{C}+\frac{\omega_{M}^{2}}{\omega_{C}}\bigg)+\mathrm{i}\,\frac{\omega_{M}^{2}\omega_{\eta}}{\omega_{C}^{2}},$ (65)
$\lambda_{3,4}$ $\approx$ $\pm\bigg(\frac{\omega_{M}^{2}}{\omega_{C}}+\frac{{\omega_{A,V}^{C}}^{2}+{\omega_{A,V}^{T}}^{2}-2\omega_{A,H}^{2}}{2\omega_{C}}\bigg)+\mathrm{i}\,\omega_{\eta}\,\bigg(1-\frac{{\omega_{A,V}^{C}}^{2}+{\omega_{A,V}^{T}}^{2}-\omega_{A,H}^{2}}{2\omega_{M}^{2}}\bigg),$ (66)
@@
$Ra_{\ell}=\frac{g\big|\alpha^{C}\beta_{y}^{C}+\alpha^{T}\beta_{y}^{T}-2\alpha^{T}\beta_{z}k_{y}/k_{z}\big|\delta^{2}}{2\varOmega\eta},$ (69)
@@
where the coefficients $D_{m}$, $G_{m}$ and $Q_{m}$ are evaluated from the initial conditions for $\hat{u}_{z}$, $\hat{\Theta}$, $\hat{\gamma}$ and their time derivatives. For instance, the initial conditions for $\hat{u}_{z}$ and its time derivatives are given by
@@
Further, $\hat{u}_{y}$ is calculated from equation (18). The spectral coefficients of $\hat{\Theta}$ and $\hat{\gamma}$ are obtained in a similar way (Appendix B).
In two-component convection, both thermal and compositional buoyancies drive the convection and are responsible for magnetic field generation. The fraction of thermal power relative to the total convective power is given by,